### v1

@@ 5 Summary and Discussion
Various types of systematic uncertainties are examined, including the GALPROP model parameters, the energy bin used to create the Fermi bubble template, and the morphology of the bubble template. Due to the strong diffuse emission from cosmic-ray interactions, the systematic uncertainty of the halo component flux becomes larger at lower photon energies. However, the trend that the halo component rises from 1 GeV, peaks at around 20 GeV, and becomes almost undetectable above 200 GeV remains unchanged. In particular, it is noteworthy that the halo-like excess is detected with a similar spectrum, even when the LAT standard background model (GIEM) is used as the zero point. Since the GIEM includes a non-template patch created from the fit residuals, this suggests that the halo excess is not an artifact dependent on the model map templates used in the baseline analysis.
@@
However, the signal from the MW halo alone does not constitute the definitive proof of dark matter annihilation. Detection of annihilation signals from other objects or regions with consistent WIMP parameters will be crucial for the final confirmation. Gamma-ray observations of dwarf galaxies in the MW halo are particularly interesting from this perspective. Given the tension with the constraints from dwarf galaxies, a modest improvement in the sensitivity of future observations should eventually lead to the detection of annihilation gamma rays. It is noteworthy that interesting excesses have already been observed in several dwarf galaxies, even if they have not been definitively detected [13]. The largest excess has been reported in Reticulum II by several studies (typically –) [73, 74, 75, 76, 13], and the WIMP mass estimated by the latest study [13] is interestingly similar to that for the MW halo excess found by the present study.
Another important avenue is the search for line gamma-ray emission. Line emissions at the photon energy equal to the WIMP mass (0.3–0.8 TeV for the channel) are expected, although the intensity will depend on the branching ratio of annihilation into two photons (or perhaps ). In this energy range, ground-based atmospheric Cherenkov telescopes are more sensitive than the Fermi LAT, and further significant improvements in sensitivity are expected shortly by the Cherenkov Telescope Array Observatory (CTAO) [77].
Dark matter search using neutrinos is also interesting. Though the current neutrino constraints [4, 78, 79, 80] are 1–2 orders of magnitude (depending on the annihilation channels) weaker than the mass and cross section suggested in this work, future improvements in the sensitivity of neutrino detectors may enable independent verification. Since the gamma-ray to neutrino flux ratio can be calculated using the Standard Model of particle physics, the neutrino flux from the Fermi halo can be predicted without uncertainty about the dark matter density profile.
@@ Acknowledgments
The author thanks Tsunefumi Mizuno for useful discussion and advice. The author was supported by the JSPS/MEXT KAKENHI Grant Number 18K03692.

### v2

@@ 5 Summary and Discussion
Various types of systematic uncertainties are examined, including the GALPROP model parameters and the morphology of the Fermi bubble template. Due to the strong diffuse emission from cosmic-ray interactions, the systematic uncertainty of the halo component flux becomes larger at lower photon energies. However, the trend that the halo component rises from 1 GeV, peaks at around 20 GeV, and becomes almost undetectable above 200 GeV remains unchanged. In particular, it is noteworthy that the halo-like excess is detected with a similar spectrum, even when the LAT standard background model (GIEM) is used as the zero point. Since the GIEM includes a non-template patch created from the fit residuals, this suggests that the halo excess is not an artifact dependent on the model map templates used in the baseline analysis.
@@
However, the signal from the MW halo alone does not constitute the definitive proof of dark matter annihilation. Detection of annihilation signals from other objects or regions with consistent WIMP parameters will be crucial for the final confirmation. Gamma-ray observations of dwarf galaxies in the MW halo are fascinating from this perspective. Given the tension between the MW halo excess and the dwarf galaxy constraints, a modest improvement in the sensitivity of future observations should eventually lead to the detection of annihilation gamma rays. It is noteworthy that interesting excesses have already been observed in several dwarf galaxies, even if they have not been definitively detected [15]. The largest excess has been reported in Reticulum II by several studies (typically –) [77, 78, 79, 80, 15], and the WIMP mass estimated by the latest study [15] is interestingly similar to that for the MW halo excess found by the present study.
Another important avenue is to search for line gamma-ray emission. Line emissions at the photon energy equal to the WIMP mass (0.3–0.8 TeV for the channel) are expected, although the intensity will depend on the branching ratio of annihilation into two photons (or perhaps ). In this energy range, ground-based atmospheric Cherenkov telescopes are more sensitive than the Fermi LAT, and further significant improvements in sensitivity are expected shortly by the Cherenkov Telescope Array Observatory (CTAO) [81].
Dark matter searches using neutrinos are also intriguing. Though the current neutrino constraints [4, 82, 83, 84] are 1–2 orders of magnitude (depending on the annihilation channels) weaker than the WIMP parameters suggested in this work, future improvements in the sensitivity of neutrino detectors may enable independent verification. Since the gamma-ray to neutrino flux ratio can be calculated using the Standard Model of particle physics, the neutrino flux from the Fermi halo can be predicted without uncertainty about the dark matter density profile.
@@ Acknowledgments
The author thanks Tsunefumi Mizuno for the useful discussion and advice. The author was supported by the JSPS/MEXT KAKENHI Grant Number 18K03692.